OpenStudy (tyrenhere):

(01.07) Solve 3(x + 1) + 6 = 33. (1 point) 10 8 8.7 −8

OpenStudy (tyrenhere):

@princeharryyy

OpenStudy (princeharryyy):

3(x + 1) + 6 = 33. 3x +3 +6 = 33 3x+9 =33 3x = 33-9 3x = 24 x=24/3 =8

OpenStudy (tyrenhere):

There more I don't understand. They have fractions

OpenStudy (tyrenhere):

Im gonna post it in another question so you can get more medals.

OpenStudy (princeharryyy):

haha. I don't care about medals. post them here only. I don't mind.

OpenStudy (tyrenhere):

ok

OpenStudy (princeharryyy):

:)

OpenStudy (tyrenhere):

Solve y/-4 + 5 = 15. (1 point) 80 −80 40 −40 y/-4 means y over negative 4 btw

OpenStudy (tyrenhere):

and you're so much help.

OpenStudy (princeharryyy):

-y/4 + 5 = 15 -y/4 = 15-5 -y = 10*4 y=-40

OpenStudy (princeharryyy):

@TyrenHere

OpenStudy (tyrenhere):

I went to go get my hotpocket

OpenStudy (tyrenhere):

and thank you so much!

OpenStudy (princeharryyy):

okay! np.

OpenStudy (princeharryyy):

Do u have any more questions or Can I go?

OpenStudy (tyrenhere):

are you still willing to help? I dont wanna be a bother

OpenStudy (princeharryyy):

it's up to u. LOL hehe

OpenStudy (tyrenhere):

Solve for x: the quantity of \[\frac{ x + 2 }{ ?4 }\] = 7. (1 point) x = 3 x = 5 x = 6 x = 10

OpenStudy (tyrenhere):

ignore that question mark xD

OpenStudy (princeharryyy):

(x+2 )/4 =7 x+2 =7*4 =28 x=28-2 =26, there is something wrong with your question check it again

OpenStudy (tyrenhere):

http://prntscr.com/cad5h9

OpenStudy (princeharryyy):

(x+4 )/2 =7 x+4 =7*2 =14 x=14-4 =10, you inverted 2 with 4 earlier

OpenStudy (tyrenhere):

Oh my bad

OpenStudy (tyrenhere):

http://prntscr.com/cad6ln 2 more

OpenStudy (princeharryyy):

1/3 (2x-8) =4 2x-8=4*3 =12 2x=12+8 x=20/2 =10

OpenStudy (tyrenhere):

http://prntscr.com/cad7qs last one

OpenStudy (princeharryyy):

D 35+x=83

OpenStudy (princeharryyy):

35 +3x = 83 (D)

OpenStudy (tyrenhere):

are you sure u dont want me to make extra questions so you can get them medals? you really deserve them

OpenStudy (princeharryyy):

Yeah. Medals aren't necessary I guess. Nit here at least. It's not olympics. LOl

OpenStudy (tyrenhere):

lmao but i honestly really appreciate your help. thanks alot

OpenStudy (princeharryyy):

Okay ! I gotta go now. BYE! Tc.